OpenStudy (anonymous):

Limit as x approaches pi/2

OpenStudy (anonymous):

\[2e^{\tan x}\div 2x-\pi \]

OpenStudy (anonymous):

Do you mean this?\[\lim_{x \rightarrow \frac{ \pi }{ 2 }} \frac{ 2e ^{\tan x} }{ 2x - \pi }\]What difficulty are you having with it? What have you tried so far?